OpenStudy (mathisfun13):

find the value of S for an infinite geometric series with r=1/6 and a1=10

jimthompson5910 (jim_thompson5910):

use the formula S = a/(1-r) in this case a = 10 r = 1/6

jimthompson5910 (jim_thompson5910):

plug those two values in and tell me what you get

OpenStudy (mathisfun13):

okay

OpenStudy (mathisfun13):

25/3

jimthompson5910 (jim_thompson5910):

no, unfortunately that is incorrect

OpenStudy (mathisfun13):

lol okayy let me tryy again

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (mathisfun13):

12

jimthompson5910 (jim_thompson5910):

good, you got it

jimthompson5910 (jim_thompson5910):

a/(1-r) = 10/(1-1/6) = 10/(5/6) = 10*(6/5) = 60/5 = 12

OpenStudy (mathisfun13):

thanksss( :

jimthompson5910 (jim_thompson5910):

yw